OpenStudy (anonymous):

48% of what number is 84?

OpenStudy (hoblos):

84 * 100/48

OpenStudy (hoblos):

=175

OpenStudy (anonymous):

x = 84 *100/48 = 8400/48 = 175

OpenStudy (anonymous):

hoblos you can't just give them the answer straight up

OpenStudy (aroub):

48/100x=84 48x=8400 x=8400/48 x=175